Calculate the final centigrade temperature to change 20litres of gas at 120°c and 1atmosphere to 25litres at atmosphere.

1
Expert's answer
2022-03-09T10:32:54-0500

The Gay-Lussac's law says

"\frac{V_2}{V_1}=\frac{T_2}{T_1}"

Hence, the final temperature

"T_2=T_1\frac{V_2}{V_1}\\
=(120+273)*\frac{25}{20}=491\:\rm K=218\:{^\circ}C"